# Maximum Independent Set; Hardness Parameter; Neutral Atoms; Rydberg Blockade; QuEra’s Aquila; Analog Quantum Computing; Adiabatic Schedules; Counterdiabatic Driving

### v2

@@ C.2 Numerical implementation
Besides the above-mentioned implementation aspects, a crucial ingredient is the correct use of the blockade subspace: when the interaction energy between Rydberg states is much larger than the Rabi strength, the exact dynamics in the full Hilbert space is well approximated by the dynamics in the subspace where only one Rydberg excitation is allowed between nearby atoms [45]. In numerical simulations, the states |$\ket{11}_{ij}$ are therefore neglected if the distance between $i$ and $j$ is smaller than a given subspace radius $R_{\text{s}}$, which results in much faster computations. The default setting in most programs is $R_{\text{s}}=R_{\text{b}}$, so that the quantum evolution is restricted to the independent sets in unit disk graphs. However, this approximation is too coarse to get accurate results, see Fig. 14: neglecting the states |$\ket{11}$ for next-nearest neighbors (extremities of a diagonal) leads to success probabilities that are overestimated by up to 50%. Therefore, it is crucial to simulate either the full Hilbert space or the blockade subspace where only nearest neighbor Rydberg states are neglected.
Optimization algorithm
@@
In this study, we focus on the probability to reach a MIS, i.e., we want to maximize the overlap between the quantum state at the end of the evolution and all MIS solutions of the graph under investigation, see Eq. (5). The set of parameters $\Theta$ to be optimized depends on the protocol, e.g., $\Theta=(\tau_{\text{i}},\tau_{\text{f}},\Delta_{\text{i}},\Delta_{\text{f}})$ for the simple linear schedules. Since the Schrödinger evolution is deterministic and that we simulate noiseless systems, any standard optimization algorithm may be applied. We choose the simplex search method [46], which turns out to be robust and very efficient in our setting. Finally, we set $t_{\text{f}}=1$ µs, which leads to success probabilities that are well distributed over the whole unit interval for the 500 graphs of the dataset.
@@
Figure 14: Numerical simulations of the Rydberg Hamiltonian for the toy graph “7DE” and the ACQC protocol defined in [28] with $t_{\text{f}}=1$ µs. The height of the bars represents the probability to find the atom $i$ in the state |$\ket{1}$ at the end of the evolution, with $i=1,\ldots,9$ from top to bottom and left to right in the graph. The MIS solution (1, 3, 7, 9) is clearly visible in this bar chart. We verify that our implementation (light colors) is perfectly equivalent to that of Bloqade (dark colors); note that, depending on the SDK, the phase has to be multiplied by minus one in Eq. (23b). We also observe that the simulations in the full Hilbert space (blue bars) match those performed in the blockade subspace of nearest neighbors (orange bars). However, the unit disk subspace approximation $R_{\text{s}}=R_{\text{b}}$ (green bars) is clearly in favor of the MIS solution and thus cannot be considered as satisfactory. This is even more noticeable when comparing the probability to reach the ground state at the end of the evolution for each blockade subspace: 0.49(4), 0.49(6), and 0.70(0) respectively.

### v3

@@ C.2.1 Blockade subspace
Besides the above-mentioned implementation aspects, a crucial ingredient is the correct use of the blockade subspace: when the interaction energy between Rydberg states is much larger than the Rabi strength, the exact dynamics in the full Hilbert space is well approximated by the dynamics in the subspace where only one Rydberg excitation is allowed between nearby atoms [39]. In numerical simulations, the states |$\ket{11}_{ij}$ are therefore ignored if the distance between $i$ and $j$ is smaller than a given subspace radius $R_{\text{s}}$, which results in much faster computations. The default setting in most programs is $R_{\text{s}}=R_{\text{b}}$, so the quantum evolution is restricted to the independent sets in unit disk graphs. However, this approximation is too coarse to get accurate results—see Fig. 15: ignoring the states |$\ket{11}$ for next-nearest neighbors (extremities of a diagonal) leads to success probabilities that are overestimated by up to 50%. Therefore, it is crucial to simulate either the full Hilbert space or the blockade subspace where only nearest-neighbor Rydberg states are ignored.
C.2.2 Optimization algorithm
@@
In this study, we focus on the probability of reaching a MIS, i.e., we want to maximize the overlap between the quantum state at the end of the evolution and all MIS solutions of the graph under investigation; see Eq. (4). The set of parameters $\Theta$ to be optimized depends on the protocol, e.g., $\Theta=(\tau_{\text{i}},\tau_{\text{f}},\Delta_{\text{i}},\Delta_{\text{f}})$ for the simple linear schedules. Since the Schrödinger evolution is deterministic and we simulate noiseless systems, any standard optimization algorithm may be applied. We choose the simplex search method [53], which turns out to be robust and very efficient in our setting. Finally, we set $t_{\text{f}}=1$ µs, which leads to success probabilities that are well distributed over the whole unit interval for the 500 graphs of the dataset.
@@
Figure 15: Numerical simulations of the Rydberg Hamiltonian for the toy graph “7DE” and the ACQC protocol defined in Ref. [32] with $t_{\text{f}}=1$ µs. The height of the bars represents the probability to find the atom $i$ in the state |$\ket{1}$ at the end of the evolution, with $i=1,\ldots,9$ from top to bottom and left to right in the graph. The MIS solution (1, 3, 7, 9) is clearly visible in this bar chart. We verify that our implementation (light colors) is perfectly equivalent to that of BLOQUADE (dark colors); note that, depending on the software development kit, the phase has to be multiplied by minus one in Eq. (23b). We also observe that the simulations in the full Hilbert space (blue bars) match those performed in the blockade subspace of nearest neighbors (orange bars). However, the unit disk subspace approximation $R_{\text{s}}=R_{\text{b}}$ (green bars) is clearly in favor of the MIS solution and thus cannot be considered as satisfactory. This is even more noticeable when comparing the probability to reach the ground state at the end of the evolution for each blockade subspace: 0.49(4), 0.49(6), and 0.70(0) respectively.